OpenStudy (sphott51):

cx+b =3(x-c) cx+b =3x-3c I don't know what to do next..

OpenStudy (evoker):

Hmm perhaps break into two equations, cx=3x and b=-3c

OpenStudy (sphott51):

how would I do that? it's solve for x

OpenStudy (evoker):

Ahh well in that case, probably need to see the original question.

OpenStudy (evoker):

or are you just asked to find the symbolic answer

OpenStudy (evoker):

If just symbollic, collect the terms with x on one side, and the ones without on the other side.

OpenStudy (evoker):

i.e subtract 3x from both sides and subtract b from both sides

OpenStudy (sphott51):

like making it into -2x +cx = -3c -b ?

OpenStudy (evoker):

yes if the first term was -3x

OpenStudy (evoker):

so -3x+cx=-3c-b

OpenStudy (evoker):

next factor x out on the left side

OpenStudy (karim728):

solving for x?

OpenStudy (welshfella):

Yes factoring out x x(c -3) = - 3c - b Now divide both sides by c - 3

OpenStudy (sphott51):

x = -3c -b / c- 3 ... ?